Carlos scored 15 more points than John in the basketball game. If the product of the two point totals is 154, how many points did each boy score?
Answer by fractalier(6550)

Let their totals be C and J.
Then C = J + 15 and
CJ = 154
Now substituting we get
J(J + 15) = 154
J^2 + 15J = 154
J^2 + 15J - 154 = 0
(J + 22)(J - 7) = 0
J = -22 or J = 7
but since you can't score negative points,
J = 7 and C = 22